# Fully tensorial approach to hypercomplex neural networks

### v3

@@ Abstract
Fully tensorial theory of hypercomplex neural networks is given. It allows neural networks to use arithmetic based on arbitrary algebras. The key point is to observe that algebra multiplication can be represented as a rank three tensor and use this tensor in every algebraic operation. This approach is attractive for neural network libraries that support effective tensorial operations. It agrees with previous implementations for four-dimensional algebras.
@@
Tensor, Hypercomplex-valued Neural Network, Hypercomplex Algebra, TensorFlow, PyTorch
@@
[inst1]organization=Faculty of Computer Science and Telecommunications, Cracow University of Technology, addressline=Warszawska 24, city=Cracow, postcode=PL31-155, country=Poland
@@ 2.1 Tensors
[…]
@@
Let us notice that the matrix $[F_{ij}]$ is implemented in TensorFlow and PyTorch library as a tensor class. The critical difference is that mathematical tensors have specific properties of transformations under the change of basis in the underlying vector spaces. The libraries implement the tensors as a multidimensional matrix of numbers. Moreover, the libraries usually do not keep the upper (contravariant) or lower (covariant) position of indices.
@@
Algebra over a field $\mathbb{F}$ is a vector space $V$ equipped with a product - a binary operation __$\_\cdot\_:V\times V\rightarrow V$ with the following properties:
@@
$\begin{array}[]{c|c|c}\cdot&e_{0}&e_{1}\\ \hline\cr e_{0}&e_{0}&e_{1}\\ \hline\cr e_{1}&e_{1}&-e_{0}\\ \end{array}$ (16)
@@ 3.1 Hypercomplex dense layer
We start with a description of the dense hypercomplex layer. It is a general-purpose layer that operates on the data with additional dimensionality/axis, which is a multiple of algebra dimensions. We assume that the input data are of dimension $b\times\underbrace{al\times in}_{n}$, where $b$ is the batch size, $al$ the algebra size, and $in$ the positive integer multiplier. The last two numbers determine the input data size $n$. The input tensor in abstract index notation is $X=X_{i_{b}Rs(i_{al},i_{in})}$, where $i_{b}$ batch index, $i_{al}$ is the algebra index, $i_{in}$ is the multiplicity index of algebra dimension. Moreover, we use learning parameters (weights/kernel) $K=K_{i_{al}i_{in}i_{u}}$, where $i_{u}$ is the index over units/neurons. The bias $b=b_{Rs(i_{al},i_{u})}$ is used if needed. Kernel and bias are usually initialized with numbers taken from specific distributions (cf. [12]).
@@
$X$, $A$, $\sigma$, bias, activation
$K$, $b$ initialized
@@
The implementation using Keras with TensorFlow library implementation and PyTorch library implementations are described in [17] and implemented in Hypercomplex Keras library [25] in the HyperDense class. As it was emphasized, for $4$-diemnsional algebras the output agrees with the results of [23].
@@ 3.2 Hypercomplex convolutional layer
The dimension of the input data $X=X_{i_{b}i_{1}\ldots i_{k}Rs(i_{al},i_{in})}$ of dimension $b\times n_{1}\times\ldots\times n_{k}\times\underbrace{al\times in}_{n}$, where $b$ is the batch size, $n_{i}$, $i\in\{1,\ldots,k\}$ the size of data sample in each dimension (e.g., for $k=2$ the raster coordinates $i_{1}$, $i_{2}$), $al$ the algebra dimension, and $in$ the positive multiplier [^3] One can see that the size of the data sample in each dimension (axes) stands right after the batch size, which is used in TensorFlow convention. For PyTorch, this multiindex is placed at the end. We will not provide an implementation for the PyTorch convention since it differs by a few permutations.. The kernel size is $al\times L_{1}\times\ldots L_{k}\times in\times F$, where: $L_{i}$, $i\in\{1,\ldots,k\}$ is the kernel dimension for each axis, and $F$ the number of filters. We therefore have the kernel $K=K_{i_{al}i_{l_{1}}\ldots i_{l_{k}}i_{in}i_{f}}$. The flag $bias$ indicates the bias $b$ usage. If it is used, it has dimension $al\times F$. Kernel and bias can be initialized by arbitrary distributions; see [12]. We use standard $k$-dimensional convolution $convkD(X,K,strides,padding)$ for convoluting algebra components, there are standard optimized convolution operations [16]. The algorithm is presented in the Algorithm 2.
@@
$X$, $A$, $\sigma$, bias, activation
$K$, $b$ initialized
@@
for $i\in\{0,\ldots,al\}$ do temp = temp + [$convkD(X,W_{i\ldots})$]
@@ 4 Tensor version of universal approximation theorem for single layer of hypercomplex perceptrons
For a Hilbert space $H$ over field $\mathbb{F}$ (e.g., $\mathbb{R}$ or $\mathbb{C}$) with inner product <_|_$<\_|\_>:X\times X\rightarrow\mathbb{F}$ every continous linear functional $\phi\in H^{*}$ there is a vector $v_{\phi}\in H$ such that
@@
When we take $H=\mathbb{R}^{N}$, for integer $N>0$, then $<x|y>:=x^{T}y=\sum_{i=1}^{n}x_{i}y_{i}$. Moreover, if $dim(H)<\infty$, then the linear functional is automatically continuous.
@@
In other words, the set of functions $S=\sum_{k=1}^{p}\alpha_{k}f(<v_{k}|x>+\Theta_{k})$ for parameters: $\alpha_{k}$, $v_{k}$ and $\Theta_{k}$ are dense in $C^{0}(X,V)$ continuous functions from $X$ to $v$ with the supremum norm over components of $V$ treated as a vector space.
@@
Then each component $\pi_{i}\circ g\in C^{0}(X,\mathbb{R})$ can be approximated by the Cybenko’s Universal Approximation Theorem [9]: there are real coefficients $a_{1}^{i},\ldots,a_{p}^{i}$, real-valued linear functionals $\phi_{1}^{i},\ldots,\phi_{p}^{i}$ on $\mathbb{R}^{nN}=V^{n}$ and some real parameters $\theta_{1}^{i},\ldots,\theta_{p}^{i}$, where the index depends on the component $i$, i.e., $p=p(i)$, such that
@@
$sup_{x\in X}||\sum_{k=1}^{p}\alpha_{k}\sigma(\phi_{k}^{i}(x)+\theta_{k}^{i})e_{i}-\sum_{k=1}^{p}\alpha_{k}f(<v_{k}^{i}|x>+\Theta_{k}^{i}(\lambda^{i}))||<\varepsilon,$ (30)

### v4

@@ Abstract
Fully tensorial theory of hypercomplex neural networks is given. It allows neural networks to use arithmetic based on arbitrary algebras. The key point is to observe that algebra multiplication can be represented as a rank three tensor and use this tensor in every algebraic operation. This approach is attractive for neural network libraries that support effective tensorial operations. It agrees with previous implementations for four-dimensional algebras. The proof of Universal Approximation Theorem for tensor formalism was given.
@@
Tensor, Hypercomplex-valued Neural Network, Hypercomplex Algebra, Universal Approximation Theorem for Tensors, Hypercomplex Keras
@@
[inst1]organization=Faculty of Computer Science and Mathematics, Cracow University of Technology, addressline=Warszawska 24, city=Cracow, postcode=PL31-155, country=Poland
@@ 2.1 Tensors
${V\times W\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{F}\scriptstyle{\otimes}$${V\otimes W\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{f}$${\mathbb{F}}$
@@
Let us notice that the matrix $F_{ij}]$ is implemented in TensorFlow and PyTorch library as a tensor class. The critical difference is that mathematical tensors have specific properties of transformations under the change of basis in the underlying vector spaces. The libraries implement the tensors as a multidimensional matrix of numbers. Moreover, the libraries usually do not keep the upper (contravariant) or lower (covariant) position of indices.
@@
Algebra over a field $\mathbb{F}$ is a vector space $V$ equipped with a product - a binary operation $\cdot\_:V\times V\rightarrow V$ with the following properties:
@@
[…]
@@ 3.1 Hypercomplex dense layer
We start with a description of the dense hypercomplex layer. It is a general-purpose layer that operates on the data with additional dimensionality/axis, which is a multiple of algebra dimensions. We assume that the input data are of dimension $b\times\underbrace{al\times in}_{n}$, where $b$ is the batch size, $al$ - the algebra size, and $in$ - the positive integer multiplier. The last two numbers determine the input data size $n$. The input tensor in abstract index notation is $X=X_{i_{b}Rs(i_{al},i_{in})}$, where $i_{b}$ - batch index, $i_{al}$ is the algebra index, $i_{in}$ is the multiplicity index of algebra dimension. Moreover, we use learning parameters (weights/kernel) $K=K_{i_{al}i_{in}i_{u}}$, where $i_{u}$ is the index over units/neurons. The bias $b=b_{Rs(i_{al},i_{u})}$ is used if needed. Kernel and bias are usually initialized with numbers taken from specific distributions (cf. [13]).
@@
$X$, AA, $\sigma$, bias, activation
$K$, $b$ - initialized
@@
The implementation using Keras with TensorFlow library implementation and PyTorch library implementations are described in [20] and implemented in Hypercomplex Keras library [28] in the HyperDense class. As it was emphasized, for 44-diemnsional algebras the output agrees with the results of [26].
@@ 3.2 Hypercomplex convolutional layer
The dimension of the input data $X=X_{i_{b}i_{1}\ldots i_{k}Rs(i_{al},i_{in})}$ of dimension $b\times n_{1}\times\ldots\times n_{k}\times\underbrace{al\times in}_{n}$, where $b$ is the batch size, $n_{i}$, $i\in\{1,\ldots,k\}$ - the size of data sample in each dimension (e.g., for $k=2$ the raster coordinates $i_{1}$, $i_{2}$), $al$ - the algebra dimension, and $in$ - the positive multiplier [^3] One can see that the size of the data sample in each dimension (axes) stands right after the batch size, which is used in TensorFlow convention. For PyTorch, this multiindex is placed at the end. We will not provide an implementation for the PyTorch convention since it differs by a few permutations.. The kernel size is $al\times L_{1}\times\ldots L_{k}\times in\times F$, where: $L_{i}$, $i\in\{1,\ldots,k\}$ is the kernel dimension for each axis, and $F$ - the number of filters. We therefore have the kernel $K=K_{i_{al}i_{l_{1}}\ldots i_{l_{k}}i_{in}i_{f}}$. The flag biasbias indicates the bias $b$ usage. If it is used, it has dimension $al\times F$. Kernel and bias can be initialized by arbitrary distributions; see [13]. We use standard $k$-dimensional convolution $convkD(X,K,strides,padding)$ for convoluting algebra components, there are standard optimized convolution operations [18]. The algorithm is presented in the Algorithm 2.
@@
$X$, AA, $\sigma$, bias, activation
$K$, $b$ - initialized
@@
for $i\in\{0,\ldots,al\}$ do temp = temp + $convkD(X,W_{i\ldots})$]
@@ 4 Tensor version of universal approximation theorem for single layer of hypercomplex perceptrons
For a Hilbert space $H$ over field $\mathbb{F}$ (e.g., $\mathbb{R}$ or $\mathbb{C}$) with inner product $|\_>:X\times X\rightarrow\mathbb{F}$ every continous linear functional $\phi\in H^{*}$ there is a vector $v_{\phi}\in H$ such that
@@
When we take $H=\mathbb{R}^{N}$, for integer $N>0$, then $x|y>:=x^{T}y=\sum_{i=1}^{n}x_{i}y_{i}$. Moreover, if $dim(H)<\infty$, then the linear functional is automatically continuous.
@@
In other words, the set of functions $S=\sum_{k=1}^{p}\alpha_{k}f(<v_{k}|x>+\Theta_{k})$ for parameters: $\alpha_{k}$, $v_{k}$ and $\Theta_{k}$ are dense in $C^{0}(X,V)$ - continuous functions from $X$ to $v$ with the supremum norm over components of $V$ treated as a vector space.
@@
Then each component $\pi_{i}\circ g\in C^{0}(X,\mathbb{R})$ can be approximated by the Cybenko’s Universal Approximation Theorem (UAP) [^4] There is a discussion about the correctness of the proof of the Cybenko’s UAP from [10] related to the usage of the Hahn-Banach theorem for measures. However, there are also other similar results that are proven using different techniques [5], [16], [19]. Therefore it seems that the theorem is firmly established even if some approaches to the proof are under considerations. [10]: there are real coefficients $a_{1}^{i},\ldots,a_{p}^{i}$, real-valued linear functionals $\phi_{1}^{i},\ldots,\phi_{p}^{i}$ on $\mathbb{R}^{nN}=V^{n}$ and some real parameters $\theta_{1}^{i},\ldots,\theta_{p}^{i}$, where the index depends on the component $i$, i.e., $p=p(i)$, such that
@@
Example An example of degenerate algebra is provided by a multiplication table/tensor $A$ that has at least one null row/column or can be brought to this form by the linear transformation of the base. Since there is a unit element in the algebra, so degenerate algebras does not have unit elements, e.g., $V=span\{e_{0}\}$ where $e_{0}e_{0}=0$ in the simplest case.
@@
[…]